one grade of tea, worth $3.90 per pound is to be mixed with another grade worth $2 per pound to make 30 pounds that will sell $2.76 per pound how much of each grade of tea must be used
Answer by mananth(16946)

component Price $ ---------------- quantity
Grade I tea 3.9 ---------------- x lb
Grade II tea 2 ---------------- 30 - x lb
Mixture 2.76% ---------------- 30

3.9 x + 2 ( 30 - x ) = 30.00 * 2.76

3.9 x + 60 - 2 x = 82.80
3.9 x - 2 x = 82.8 - 60
1.9 x = 22.8
/ 1.9
x = 12 lb $3.90 Grade I tea
18 lb $2.00 Grade II tea